# Chevalley formulae for the motivic Chern classes of Schubert cells and for the stable envelopes

### v3

@@ Abstract.
L. C. Mihalcea was supported in part by the NSF grant DMS-2152294 and a Simons Collaboration Grant. H. Naruse was supported in part by JSPS KAKENHI Grant Number 16H03921.
@@ 1. Introduction
Let $G$ be a complex, semisimple, Lie group and $T\subset B\subset P\subset G$ be a parabolic subgroup containing a Borel subgroup and the (standard) maximal torus. Let $W$ be the Weyl group determined by $(G,T)$. In the study of cohomology and K-theory rings of (generalized) flag manifolds $G/P$, the Chevalley formula expresses the multiplication of a Schubert class by the class of line bundle, or, equivalently, a Schubert divisor in $G/P$. If one works equivariantly, this formula determines completely the multiplication in the equivariant K ring. In this paper we prove a Chevalley formula for the coefficients $C_{u,\lambda}^{w}(y)\in K_{T}(pt)[y]$ arising in the multiplication
@@
The motivic Chern classes of Schubert cells generalize well studied classes from Schubert calculus. If $y=0$, the motivic class $MC_{y}(X(w)^{\circ})$ is equal to the class of the ideal sheaf $[{\mathcal{O}}_{X(w)}(-\partial X(w))]$ of the boundary $\partial X(w)=X(w)\setminus X(w)^{\circ}$, where $X(w)=\overline{X(w)^{\circ}}$ is the Schubert variety. If $y=-1$, then $MC_{y}(X(w)^{\circ})$ is equal to the class of the unique $T$-fixed point in $X(w)^{\circ}$; see [AMSS22]. The Poincaré duals of the classes $MC_{y}(X(w)^{\circ})$, the Segre motivic classes [AMSS19, MSA22], specialize when $y=0$ to the Grothendieck classes of the structure sheaves of the opposite Schubert varieties. Our Chevalley formula (1), and its analogous formula for the Segre motivic classes, specialize to known Chevalley formulae for K-theoretic Schubert classes and ideal sheaves from [GR04, LP07].
@@
We now give a rough idea on the proof of Theorem 1.1. The key connection between the Chevalley formula in the Hecke algebra to motivic Chern classes, proved in [MNS22b], and ultimately based on results from [AMSS19], is that the motivic Chern classes are recursively obtained by certain left Demazure–Lusztig operators $\mathcal{T}_{w}^{L}$ acting on $K_{T}(G/B)[y]$:
@@
A remarkable property of the motivic Chern classes of Schubert cells, proved in [AMSS19, FRW21], is that they are equivalent to the K-theoretic version of Maulik and Okounkov’s stable envelopes, see [MO19, AO21]. The stable envelopes are elements in the $T\times{\mathbb{C}}^{*}$-equivariant K-theory of the cotangent bundle, $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$. In this context, the formal variable $y$ may be identified to the (inverse) of the character given by the ${\mathbb{C}}^{*}$ fibre dilation on the cotangent bundle. If $\iota:G/B\hookrightarrow T^{*}(G/B)$ is the inclusion of the zero section, then $\iota^{*}(\operatorname{stab}(w))$ is a multiple of the motivic Chern class of the (opposite) Schubert cell for $w$, where $\operatorname{stab}$ is a stable envelope, appropriately normalized.
The stable envelopes depend on three parameters: a chamber, a polarization, and a slope, and the precise normalizations are essential for this paper. A variation in the chamber results in conjugating by the Borel subgroup [AMSS19], and it is encoded in the left Weyl group action [MNS22b] and certain $R$-matrix operators [RTV15, RTV19]. Varying the polarization, or the slope, results in the multiplication of $\operatorname{stab}(w)$ by a line bundle $\mathcal{L}_{\lambda}$ pulled back from $G/B$; cf. [AMSS19, Oko17], see also §7 below. In particular, the coefficients $C_{v,\lambda}^{w}(y)$ from (1) give ‘wall-crossing’ formulae, recording the change of stable envelopes when its defining parameters are varied. While these wall crossing formulae have been worked out in [Oko17, SZZ20, SZZ21] (see also [KW23]), in §7 we revisit some of these from the point of view of Theorem 1.1. In particular, we utilize the Chevalley formula to give an explicit combinatorial rule relating the stable envelope for the fundamental alcove $\operatorname{A_{\circ}}$ to the one corresponding to any translation $\operatorname{A_{\circ}}+\lambda$; see Theorem 7.8.
@@
Finally, in an Appendix we obtain an analogue of the Chevalley formula (1) for the homological analogue of the motivic Chern classes, the Chern–Schwartz–MacPherson classes. While this formula may be obtained by a specialization argument as in [AMSS22], the degenerate affine Hecke algebra is much simpler in this case, and a direct proof of the Chevalley formula may be obtained rather quickly.
@@
For any $w\in W$, let $X(w)^{\circ}:=BwB/B\subset G/B$ and $Y(w)^{\circ}:=B^{-}wB/B\subset G/B$ be Schubert cells, where $B^{-}$ is the opposite Borel subgroup. Let $X(w):=\overline{X(w)^{\circ}}$ and $Y(w):=\overline{Y(w)^{\circ}}$ be the Schubert varieties, respectively. Let $P(\supseteq B)$ be a parabolic subgroup with simple roots $\Sigma_{P}\subset\Sigma$. Let $R_{P}^{+}$ denote the positive roots spanned by $\Sigma_{P}$. Let $W_{P}$ be the Weyl group generated by the simple reflections $s_{\alpha}$, $\alpha\in\Sigma_{P}$. Let $W^{P}\simeq W/W_{P}$ denote the set of minimal length representatives. For any $w\in W^{P}$, let $X(wW_{P})^{\circ}:=BwP/P\subset G/P$ (resp. $Y(wW_{P})^{\circ}:=B^{-}wP/P\subset G/P$) denote the Schubert cell with closure $X(wW_{P})$ (resp. $Y(wW_{P})$). Let $X^{*}(T)_{P}:=\{\lambda\in X^{*}(T)\mid\langle\lambda,\gamma^{\vee}\rangle=0\text{ for all }\gamma\in R^{+}_{P}\}$ be the set of integral weights which vanish on $(R^{+}_{P})^{\vee}$. For any $\lambda\in X^{*}(T)_{P}$, we still use ${\mathcal{L}}_{\lambda}$ to denote the line bundle $G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}(G/P)$, which has fiber over $1.P$ the $T$-module of weight $\lambda$.
@@ 2.2.2. Alcove walk algebra
Let $G=\operatorname{SL}(2,{\mathbb{C}})$. We use the same notation as in Example 2.1. We check the above theorem for $w=s_{1}$ and $\lambda=\omega$. From the alcove picture in Example 2.1, $T_{s_{1}}^{-1}$ is represented by the minimal length walk $q^{-1/2}c_{1}^{-1}$, while $X^{\omega}$ is represented by the walk $gc_{1}^{+}=c_{0}^{+}g$. Thus, the sum in the right hand side of the Theorem is over the alcove walks $c_{1}^{-}c_{0}^{-}g$ and $c_{1}^{-}f_{0}^{+}g$, which end at the alcoves $t_{-\omega}s_{1}\operatorname{A_{\circ}}$ and $t_{-\omega}\operatorname{A_{\circ}}$, respectively. (Note that $c_{1}^{-}c_{0}^{+}g$, which represents $T_{s_{1}}^{-1}X^{\omega}$, is not an alcove walk.) Therefore, the identity in the theorem is
@@
Let $h_{i}=H_{\beta_{i},k_{i}}$, for $1\leq i\leq s$ be the sequence of hyperplanes defined by the walls of alcoves crossed by $p_{\lambda}$, with $\beta_{i}\in R^{\epsilon_{i}}$ and $k_{i}\in{\mathbb{Z}}$. Then there is a bijection between the following two sets:
@@ 4.1. Definition of the Motivic Chern classes
There is also an equivariant version of the motivic Chern class transformation, which uses equivariant varieties and morphisms, and has values in the suitable equivariant K-theory group. Its definition was given in [AMSS19, FRW21], following closely the approach of [BSY10].
Assume $X$ is smooth and there is a torus $T$ acting on $X$. Let $K_{T}(X)$ denote the equivariant K-theory of $X$, see [CG09]. If $X$ is a point, $K_{T}(pt)=K^{0}(\operatorname{Rep}(T))={\mathbb{Z}}[T]$. For any ${\mathcal{F}}\in K_{T}(X)$, let
@@
[AMSS19, FRW21] Let $X$ be a quasi-projective, non-singular, complex algebraic variety with an action of the torus $T$. There exists a unique natural transformation $MC_{y}:G_{0}^{T}(var/X)\to K_{T}(X)[y]$ satisfying the following properties:
@@ 4.3. Partial flag variety case
For any $w\in W$, let $\ell(wW_{P})$ denote the length of the minimal length representative in $wW_{P}$. Let $\pi:G/B\to G/P$ be the natural projection. It is proved in [AMSS19, Remeark 5.7] that
@@ 5.1. Chevalley coefficients
(30) ${\mathcal{T}}^{L}_{w}e^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}{\mathcal{T}}^{L}_{u}\in\operatorname{End}_{{\mathbb{C}}}K_{T}(G/B)[y].$
@@ 5.5. Parabolic case
In a paper in preparation we use this theorem to give a combinatorial Chevalley formula for minuscule flag varieties and a $K$-theoretic generalization of Nakada’s colored hook formula. We also recently learned about work by Neil Fan, Peter Guo, Changjian Su, and Rui Xiong who proved Pieri formulae for the motivic Chern classes on Grassmannians.
@@ 6. Dualities of Chevalley coefficients
The star duality *** acts on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(pt)[y^{\pm 1}]$ by sending a vector bundle to its dual, and leaving $y^{i}$ unchanged. Consider the composition
@@
Then we have the following result by combining Theorem 9.1(a) and Remark 4.7 of [AMSS22].
@@
Applying the star duality functor *** to Equation 27, then use Equation 34 to get
@@
    $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}.$
@@ 7. K-theoretic stable envelopes for $T^{*}(G/B)$
In this section we apply the Chevalley formula for motivic classes to calculate the transformation of stable envelopes in $T^{*}(G/B)$ under the change of arbitrary alcoves. For alcoves adjacent to the fundamental alcove, a formula for this transformation was obtained in [SZZ20, Theorem 5.4], see also [KW23].
@@ 7.1. Definition of the stable envelopes
Strictly speaking, $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\in K_{T\times{\mathbb{C}}^{*}}(G/B)$ denotes the image of $1\in K_{T\times{\mathbb{C}}^{*}}(e_{w})$ under the map $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}$. From the definition, it is immediate to see that $\{\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\mid w\in W\}$ forms a basis for the localized equivariant K-theory $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))_{\operatorname{loc}}$, called the stable basis. Explicit combinatorial formulae and recursions for the localizations of stable envelopes in $T^{*}(G/B)$ may be found in [AMSS19, §8.3].

### v4

@@ Abstract.
L. C. Mihalcea was supported in part by the NSF grant DMS-2152294 and a Simons Collaboration Grant. H. Naruse was supported in part by JSPS KAKENHI Grant Number 16H03921. C. Su is supported by the National Key R&D Program of China (No. 2024YFA1014700).
@@ 1. Introduction
Let $G$ be a complex, semisimple, Lie group and $T\subset B\subset P\subset G$ be a parabolic subgroup containing a Borel subgroup and the (standard) maximal torus. Let $W$ be the Weyl group determined by $(G,T)$. In the study of cohomology and K-theory rings of (generalized) flag manifolds $G/P$, the Chevalley formula expresses the multiplication of a Schubert class by the class of a line bundle, or a Schubert divisor in $G/P$. If one works equivariantly, this formula determines completely the multiplication in the equivariant K ring. In this paper we prove a Chevalley formula for the coefficients $C_{u,\lambda}^{w}(y)\in K_{T}(pt)[y]$ arising in the multiplication
@@
The motivic Chern classes of Schubert cells generalize well studied classes from Schubert calculus. If $y=0$, the motivic class $MC_{y}(X(w)^{\circ})$ is equal to the class of the ideal sheaf $[{\mathcal{O}}_{X(w)}(-\partial X(w))]$ of the boundary $\partial X(w)=X(w)\setminus X(w)^{\circ}$, where $X(w)=\overline{X(w)^{\circ}}$ is the Schubert variety. If $y=-1$, then $MC_{y}(X(w)^{\circ})$ is equal to the class of the unique $T$-fixed point in $X(w)^{\circ}$; see [AMSS24b]. The Poincaré duals of the classes $MC_{y}(X(w)^{\circ})$, the Segre motivic classes [AMSS24a, MS22], specialize when $y=0$ to the Grothendieck classes of the structure sheaves of the opposite Schubert varieties. Our Chevalley formula (1), and its analogous formula for the Segre motivic classes, specialize to known Chevalley formulae for K-theoretic Schubert classes and ideal sheaves from [GR04, LP07].
@@
As mentioned above, our Chevalley formula generalizes to the motivic situation the classical Chevalley multiplication in $K_{T}(G/B)$. It also generalizes the Chevalley multiplication by (equivariant) Chern–Schwartz–MacPherson classes of Schubert cells from [AMSS23]; a short, self-contained, proof of this is added in an Appendix.
@@
We now give a rough idea on the proof of Theorem 1.1. The key connection between the Chevalley formula in the Hecke algebra to motivic Chern classes, proved in [MNS22b], and ultimately based on results from [AMSS24a], is that the motivic Chern classes are recursively obtained by certain left Demazure–Lusztig operators $\mathcal{T}_{w}^{L}$ acting on $K_{T}(G/B)[y]$:
@@
A remarkable property of the motivic Chern classes of Schubert cells, proved in [AMSS24a, FRW21], is that they are equivalent to the K-theoretic version of Maulik and Okounkov’s stable envelopes, see [MO19, AO21]. The stable envelopes are elements in the $T\times{\mathbb{C}}^{*}$-equivariant K-theory of the cotangent bundle, $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$. In this context, the formal variable $y$ may be identified to the (inverse) of the character given by the ${\mathbb{C}}^{*}$ fibre dilation on the cotangent bundle. If $\iota:G/B\hookrightarrow T^{*}(G/B)$ is the inclusion of the zero section, then $\iota^{*}(\operatorname{stab}(w))$ is a multiple of the motivic Chern class of the (opposite) Schubert cell for $w$, where $\operatorname{stab}$ is a stable envelope, appropriately normalized.
The stable envelopes depend on three parameters: a chamber, a polarization, and a slope, and the precise normalizations are essential for this paper. A variation in the chamber results in conjugating by the Borel subgroup [AMSS24a], and it is encoded in the left Weyl group action [MNS22b] and certain $R$-matrix operators [RTV15, RTV19]. Varying the polarization, or the slope, results in the multiplication of $\operatorname{stab}(w)$ by a line bundle $\mathcal{L}_{\lambda}$ pulled back from $G/B$; cf. [AMSS24a, Oko17], see also §7 below. In particular, the coefficients $C_{v,\lambda}^{w}(y)$ from (1) give ‘wall-crossing’ formulae, recording the change of stable envelopes when its defining parameters are varied. While these wall crossing formulae have been worked out in [Oko17, SZZ20, SZZ21] (see also [KW25]), in §7 we revisit some of these from the point of view of Theorem 1.1. In particular, we utilize the Chevalley formula to give an explicit combinatorial rule relating the stable envelope for the fundamental alcove $\operatorname{A_{\circ}}$ to the one corresponding to any translation $\operatorname{A_{\circ}}+\lambda$; see Theorem 7.8.
@@
Finally, in an Appendix we obtain an analogue of the Chevalley formula (1) for the homological analogue of the motivic Chern classes, the Chern–Schwartz–MacPherson classes. While this formula may be obtained by a specialization argument as in [AMSS24b], the degenerate affine Hecke algebra is much simpler in this case, and a direct proof of the Chevalley formula may be obtained rather quickly.
@@
For any $w\in W$, let $X(w)^{\circ}:=BwB/B\subset G/B$ and $Y(w)^{\circ}:=B^{-}wB/B\subset G/B$ be Schubert cells, where $B^{-}$ is the opposite Borel subgroup. Let $X(w):=\overline{X(w)^{\circ}}$ and $Y(w):=\overline{Y(w)^{\circ}}$ be the Schubert varieties, respectively. Let $P(\supseteq B)$ be a parabolic subgroup with simple roots $\Sigma_{P}\subset\Sigma$. Let $R_{P}^{+}$ denote the positive roots spanned by $\Sigma_{P}$. Let $W_{P}$ be the Weyl group generated by the simple reflections $s_{\alpha}$, $\alpha\in\Sigma_{P}$. Let $W^{P}\simeq W/W_{P}$ denote the set of minimal length representatives. For any $w\in W^{P}$, let $X(wW_{P})^{\circ}:=BwP/P\subset G/P$ (resp. $Y(wW_{P})^{\circ}:=B^{-}wP/P\subset G/P$) denote the Schubert cell with closure $X(wW_{P})$ (resp. $Y(wW_{P})$). Let $X^{*}(T)_{P}:=\{\lambda\in X^{*}(T)\mid\langle\lambda,\gamma^{\vee}\rangle=0\text{ for all }\gamma\in R^{+}_{P}\}$ be the set of integral weights which vanish on $(R^{+}_{P})^{\vee}$. For any $\lambda\in X^{*}(T)_{P}$, we still use ${\mathcal{L}}_{\lambda}$ to denote the line bundle $G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}(G/P)$, which has fiber over $1.P$ the one dimensional $T$-module of weight $\lambda$.
@@ 2.2.1. Alcoves
We consider the example of the root system of type $A_{1}$. Let $\alpha$ be the positive root, and $\omega=\alpha/2$ be the fundamental weight. The weight lattice $X^{*}(T)={\mathbb{Z}}\omega$, the root lattice $Q={\mathbb{Z}}\alpha$, and the finite Weyl group $W=\{id,s_{1}=s_{\alpha}\}$. The affine Weyl group $W_{\operatorname{aff}}=Q\rtimes W$ has Coxeter generators $s_{1}$ and $s_{0}=t_{\alpha}s_{1}$. The subgroup of length zero elements in $W^{\textit{ext}}_{\operatorname{aff}}$ is $\Omega=\{id,g=t_{\omega}s_{1}\}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$.
In the following picture for $\Omega\times\mathfrak{t}_{\mathbb{R}}^{*}$, the lower sheet is the identity sheet, while the upper sheet is the sheet $g\times\mathfrak{t}_{\mathbb{R}}^{*}$. Each alcove $w\operatorname{A_{\circ}}$ is labeled by the corresponding $w\in W^{\textit{ext}}_{\operatorname{aff}}$, both in the Coxeter presentation and the translation presentation. In the lower sheet, the walls $H_{\alpha,n}$ are labeled by $1$ if $n$ is even, and 00 if $n$ is odd. On the upper sheet, the labelings are in the opposite way.
@@ 2.2.2. Alcove walk algebra
Let $G=\operatorname{SL}(2,{\mathbb{C}})$. We use the same notation as in Example 2.1. We check the above theorem for $w=s_{1}$ and $\lambda=\omega$. From the alcove picture in Example 2.1, $T_{s_{1}}^{-1}$ is represented by the minimal length walk $q^{-1/2}c_{1}^{-}$, while $X^{\omega}$ is represented by the walk $gc_{1}^{+}=c_{0}^{+}g$. Thus, the sum in the right hand side of the Theorem is over the alcove walks $c_{1}^{-}c_{0}^{-}g$ and $c_{1}^{-}f_{0}^{+}g$, which end at the alcoves $t_{-\omega}s_{1}\operatorname{A_{\circ}}$ and $t_{-\omega}\operatorname{A_{\circ}}$, respectively. (Note that $c_{1}^{-}c_{0}^{+}g$, which represents $T_{s_{1}}^{-1}X^{\omega}$, is not an alcove walk.) Therefore, the identity in the theorem is
@@
Let $h_{i}=H_{\beta_{i},k_{i}}$, for $1\leq i\leq s$ be the sequence of hyperplanes defined by the walls of alcoves crossed by $p_{\lambda}$, with $\beta_{i}\in R^{\epsilon_{i}}$ and $k_{i}\in{\mathbb{Z}}$.
Then there is a bijection between the following two sets:
@@ 4.1. Definition of the Motivic Chern classes
There is also an equivariant version of the motivic Chern class transformation, which uses equivariant varieties and morphisms, and has values in the suitable equivariant K-theory group. Its definition was given in [AMSS24a, FRW21], following closely the approach of [BSY10].
Assume $X$ is smooth and there is a torus $T$ acting on $X$. Let $K_{T}(X)$ denote the equivariant K-theory of $X$, see [CG10]. If $X$ is a point, $K_{T}(pt)=K^{0}(\operatorname{Rep}(T))={\mathbb{Z}}[T]$. For any ${\mathcal{F}}\in K_{T}(X)$, let
@@
[AMSS24a, FRW21] Let $X$ be a quasi-projective, non-singular, complex algebraic variety with an action of the torus $T$. There exists a unique natural transformation $MC_{y}:G_{0}^{T}(var/X)\to K_{T}(X)[y]$ satisfying the following properties:
@@ 4.3. Partial flag variety case
For any $w\in W$, let $\ell(wW_{P})$ denote the length of the minimal length representative in $wW_{P}$. Let $\pi:G/B\to G/P$ be the natural projection. It is proved in [AMSS24a, Remeark 5.7] that
@@ 5.1. Chevalley coefficients
${\mathcal{T}}^{L}_{w}e^{-\lambda}=\sum_{\mu\in X^{*}(T),u\in W}y^{\ell(w)-\ell(u)}e^{-\mu}c_{u,\mu}^{w,\lambda}|_{q=-y}{\mathcal{T}}^{L}_{u}\in\operatorname{End}_{{\mathbb{C}}}K_{T}(G/B)[y].$
@@ 5.5. Parabolic case
In a paper in preparation we use this theorem to give a combinatorial Chevalley formula for minuscule flag varieties and a $K$-theoretic generalization of Nakada’s colored hook formula. See also [FGSX24] for a Pieri formula for the motivic Chern classes of Schubert cells in the equivariant K-theory of Grassmannians.
@@ 6. Dualities of Chevalley coefficients
The star duality * acts on $K_{T}(G/B)[y^{\pm 1}]$ and $K_{T}(pt)[y^{\pm 1}]$ by sending a vector bundle to its dual, and leaving $y^{i}$ unchanged. Consider the composition
@@
Then we have the following result by combining Theorem 9.1(a) and Remark 4.7 of [AMSS24b].
@@
Applying the star duality functor * to Equation 27, then use Equation 32 to get
@@
    $(C^{w}_{u,\lambda})|_{y\to y^{-1}}\times y^{\ell(w)-\ell(u)}=C^{w}_{u,\lambda}\/.$
@@ 7. K-theoretic stable envelopes for $T^{*}(G/B)$
In this section we apply the Chevalley formula for motivic classes to calculate the transformation of stable envelopes in $T^{*}(G/B)$ under the change of arbitrary alcoves. For alcoves adjacent to the fundamental alcove, a formula for this transformation was obtained in [SZZ20, Theorem 5.4], see also [KW25].
@@ 7.1. Definition of the stable envelopes
Strictly speaking, $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\in K_{T\times{\mathbb{C}}^{*}}(G/B)$ denotes the image of $1\in K_{T\times{\mathbb{C}}^{*}}(e_{w})$ under the map $\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}$. From the definition, it is immediate to see that $\{\operatorname{stab}_{\mathfrak{C},T^{\frac{1}{2}},A}(w)\mid w\in W\}$ forms a basis for the localized equivariant K-theory $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))_{\operatorname{loc}}$, called the stable basis. Explicit combinatorial formulae and recursions for the localizations of stable envelopes in $T^{*}(G/B)$ may be found in [AMSS24a, §8.3].